OpenStudy (anonymous):

(4x+1)(2x^2-2) how do you do foil for this ? @barbie15

OpenStudy (mathstudent55):

F: first times first: \(4x \times 2x^2 = 8x^3\) O: outside terms: \(4x \times (-2) = -8x\) I: inside terms: \(1 \times 2x^2 = 2x^2\) L: last times last: \(1 \times (-2) = -2\) Add all terms together: \(8x^3 - 8x + 2x^2 - 2\) Now if you'd like you can arrange the terms in descending order of degree.